# Deparametrization and Quantization of Scalar-Tensor Gravity and Its Cosmological Model

### v1

@@ III.1 Deparametrization of the Scalar-Tensor Theory
which satisfies $t(\phi,A,E)(x)\geq 0$ on the constraint surface. Hence the Hamiltonian constraint can be reformulated as
$C(x)=\pi(x)-(-\frac{\tilde{K}_{a}^{i}E_{i}^{a}}{\phi}\pm\sqrt{t(\phi,A,E)})$ (III.3)
The deparametrized system is composed of the Gaussian constraint (II.26), the diffeomorphism constraint (II.28) and the new Hamiltonian constraint (III.3), which is equivalent to the original constraint system. It remains a first-class constraint system[4].
@@
$\pi\!=\!\pm\!\sqrt{\frac{2p_{ab}p^{ab}}{\phi^{2}}\!-\!\frac{hR}{2}\!-\!\frac{3h}{4\phi^{2}}\!(D_{a}\phi)\!D^{a}\phi\!+\!\frac{h}{\phi}D_{a}D^{a}\phi\!+\!\frac{h\xi(\phi)}{\phi}}$ (III.4)
@@
$-\frac{3h}{4\phi^{2}}(D_{a}\phi)D^{a}\phi+\frac{h}{\phi}D_{a}D^{a}\phi+\frac{h\xi(\phi)}{\phi}$ (III.6)
@@
$C^{\prime}(x)=\pi(x)-(\pm\sqrt{t^{\prime}(\phi,A,E)})$ (III.7)
The deparametrized system is composed of the Gaussian constraint (II.26), the diffeomorphism constraint (II.28), the conformal constraint (II.32) and the new Hamiltonian constraint E(III.7), which is equivalent to the original constraint system.
@@ III.2 Loop Quantization of the Scalar-Tensor Theory
$Cyl^{geo}:=Span\{|\beta,\vec{j},\vec{\iota}\rangle\}$ (III.9)
@@
$\left\langle\psi\mid\psi^{\prime}\right\rangle=\int dg_{1}\ldots dg_{n^{\prime\prime}}\bar{\psi}\left(g_{1},\ldots,g_{n^{\prime\prime}}\right)\psi^{\prime}\left(g_{1},\ldots,g_{n^{\prime\prime}}\right)$ (III.11)
@@
$E(\mathrm{S},\mathrm{g}):=\int_{\mathrm{S}}\epsilon_{abc}E_{i}^{a}\mathrm{g}^{i}$ (III.12)
@@
$\phi:\Sigma\rightarrow\mathbb{R}$ (III.14)
@@
$\langle\phi\mid\phi\rangle=1,\quad\mathrm{and}\quad\left\langle\phi\mid\phi^{\prime}\right\rangle=0\quad\mathrm{whenever}\quad\phi\neq\phi^{\prime}$ (III.16)
@@
$|\phi\rangle[\pi]=\exp\left(-i\int\mathrm{d}^{3}x\phi(x)\pi(x)\right)$ (III.17)
@@
$\hat{\phi}(x)\cdot|\phi\rangle=\phi(x)|\phi\rangle$ (III.18)
@@
$\exp\left(-i\int\mathrm{d}^{3}x\mathrm{f}(x)\hat{\pi}(x)\right)\cdot|\phi\rangle=|\phi+\mathrm{f}\rangle$ (III.19)
@@
$\widehat{\partial_{a}\ldots\partial_{b}\phi}(x)\cdot|\phi\rangle=\partial_{a}\ldots\partial_{b}\phi(x)|\phi\rangle$ (III.20)
@@
$\mathcal{H}_{kin}^{t}=\mathcal{H}_{kin}^{geo}\otimes\mathcal{H}_{kin}^{sf}$ (III.21)
@@
$\hat{U}_{f}\cdot|\beta,\vec{j},\vec{\iota}\rangle\otimes|\phi\rangle=|f(\beta),\vec{j}\circ f^{-1},\vec{\iota}\circ f^{-1}\rangle\otimes|\phi\circ f^{-1}\rangle$ (III.22)
@@
$\hat{P}_{Diff_{\beta}}\cdot|\Psi_{\beta}\rangle:=\frac{1}{n_{\beta}}\sum_{g\in GS_{\beta}}\hat{U}_{g}\cdot|\Psi_{\beta}\rangle$ (III.23)
@@
$\eta\left(\Psi_{\beta}\right)\left[\Phi_{\beta^{\prime}}\right]:=\sum_{f\in Diff/Diff_{\beta}}\left\langle\hat{U}_{f}\hat{P}_{Diff_{\beta}}\Psi_{\beta}\mid\Phi_{\beta^{\prime}}\right\rangle_{kin}$ (III.24)
@@
$\left\langle\eta\left(\Psi_{\beta}\right)\mid\eta\left(\Phi_{\beta^{\prime}}\right)\right\rangle_{Diff}:=\eta\left(\Psi_{\beta}\right)\left[\Phi_{\beta^{\prime}}\right]$ (III.25)
The diffeomorphism invariant Hilbert space $\mathcal{H}_{Diff}$ is defined by the completion of $\eta[Cyl_{G}]$ with respect to the inner product given by Eq. (III.25).
@@ III.3 Quantum Dynamics of the Scalar-Tensor Theory
In the sertor of $\omega(\phi)\neq-\frac{3}{2}$, to quantizethe deparametrized Hamiltonian constraint (III.3), we devide its smearing version into the following three terms:
$C_{1}(N)=\int d^{3}xN(x)\pi(x),$ (III.26)
@@
$\begin{split}\widehat{C}_{1}(N)=&\frac{1}{2i\lambda_{0}}\biggl(\exp\left[i\lambda_{0}\int d^{3}xN(x)\hat{\pi}(x)\right]\\ &-\exp\left[-i\lambda_{0}\int d^{3}xN(x)\hat{\pi}(x)\right]\biggr)\end{split}$ (III.27)
@@
$\times$ $\exp\left(-i\lambda_{0}\int d^{3}y\pi(y)\right)$ (III.28)
@@
$=\frac{-2sgn(\phi(x))}{\lambda_{0}}\left(\sqrt{|\phi(x)|}-\sqrt{|\phi(x)+\lambda_{0}|}\right)|\phi\rangle$ (III.29)
@@
$E_{i}^{a}$ $=\frac{1}{2}\varepsilon^{abc}\varepsilon_{ijk}e_{b}^{j}e_{c}^{k}\operatorname{sgn}\left(\operatorname{det}\left(e_{a}^{i}\right)\right)$ (III.31)
$\operatorname{sgn}\left(\operatorname{det}\left(e_{a}^{i}\right)\right)$ $=\frac{1}{6}\frac{1}{\sqrt{h}}\varepsilon_{ijk}\varepsilon^{abc}e_{a}^{i}e_{b}^{j}e_{c}^{k}$ (III.32)
$e_{a}^{i}$ $=\frac{2}{\gamma}\left\{A_{a}^{i},V\right\}$ (III.33)
$\tilde{K}_{a}^{i}$ $=\gamma^{-\frac{5}{2}}\left\{A_{a}^{i},\left\{H^{E}(1),V\right\}\right\}$ (III.34)
@@
$H^{E}(1)=\frac{1}{2}\int d^{3}xF_{ab}^{j}\frac{\varepsilon_{jkl}E_{k}^{a}E_{l}^{b}}{\sqrt{h}}$ (III.35)
@@
$\times\left\{A_{f}^{n}(y),V_{(y,\epsilon)}^{\frac{2}{3}}\right\}$ (III.36)
@@
$h_{s_{I}}$ $\approx 1-\delta\dot{s}_{I}^{a}A_{a}^{i}\tau_{i}$ (III.37)
$\int d^{3}x\varepsilon^{abc}$ $\approx\sum_{\Delta}\frac{1}{6}\delta^{3}\epsilon\left(s_{I}s_{J}s_{K}\right)\varepsilon^{IJK}\dot{s}_{I}^{a}\dot{s}_{J}^{b}\dot{s}_{K}^{c}$ (III.38)
@@
$\times h_{s_{M^{\prime}}\!\left(\!\Delta^{\prime}\!\right)\!}\!\left\{\!h_{s_{M^{\prime}}\!\left(\!\Delta^{\prime}\!\right)\!}^{-1},V_{\!\left(\!v\left(\Delta^{\prime}\!\right)\!,\epsilon\!\right)\!}^{\frac{2}{3}}\right\}h_{s_{N^{\prime}}\!\left(\!\Delta^{\prime}\!\right)\!}\!\left\{\!h_{s_{N^{\prime}}\!\left(\!\Delta^{\prime}\!\right)\!}^{-1},V_{\!\left(\!v\left(\Delta^{\prime}\!\right)\!,\epsilon\!\right)\!}^{\frac{2}{3}}\!\right\}\!\Big)$ (III.39)
@@
$\quad\left.\cdot\hat{h}_{s_{N}\left(\Delta^{\prime}\right)}\left[\hat{h}_{s_{N}\left(\Delta^{\prime}\right)}^{-1},\hat{V}_{(v(\Delta^{\prime}),\epsilon)}^{\frac{2}{3}}\right]\right)\cdot|\beta,\vec{j},\vec{\iota}\rangle\otimes|\phi\rangle$ (III.40)
@@
$C_{3}(N)=\mp\lim_{\rho\rightarrow 0}\sum_{\square^{\rho}}\operatorname{sgn}\left(N_{\square^{\rho}}\right)\sqrt{T_{\square^{\rho}}}$ (III.41)
@@
$-\left(\int_{\square^{\rho}}d^{3}xN(x)\sqrt{\frac{(3+2\omega(\phi))\xi(\phi)}{\phi}}\sqrt{h}\right)^{2}$ (III.42)
@@
$H^{L}=\frac{1}{2}\frac{\varepsilon_{jmn}\varepsilon_{jkl}\tilde{K}_{a}^{m}\tilde{K}_{b}^{n}E_{k}^{a}E_{l}^{b}}{\sqrt{h}}$ (III.43)
@@
$\cdot\sum_{v\in\beta(v)\cap\square^{\rho}}N(v)(3+2\bar{\omega}(v))\hat{H}^{E}_{v}\cdot|\beta,\vec{j},\vec{\iota}\rangle\otimes|\phi\rangle$ (III.44)
@@
$\cdot\!\sum_{v\in\beta(v)\cap\square^{\rho}}\!N(v)(3\!+\!2\bar{\omega}(v))(\gamma^{2}\!+\!D^{2}(v))\hat{H}^{L}_{v}\cdot|\beta,\vec{j},\vec{\iota}\rangle\otimes|\phi\rangle$ (III.45)
@@
$\cdot\sum_{v(\Delta)=v}\hat{V}_{(v(\Delta),\epsilon)}\cdot|\beta,\vec{j},\vec{\iota}\rangle\otimes|\phi\rangle$ (III.46)
For the term $T_{\square^{\rho}}^{3}$, we use the point-splitting scheme and apply Eq. (III.31)–Eq. (III.34) to express it as
@@
$\times\left\{A_{e}^{m}(y),V_{(y,\epsilon)}^{\frac{3}{4}}\right\}\left\{A_{f}^{n}(y),V_{\left(y,\epsilon\right)}^{\frac{3}{4}}\right\}$ (III.47)
Then, by Eq. (III.37)–Eq. (III.38), it can be regularized as
@@
$\times\Big(\phi(f(s_{L}(\Delta^{\prime})))-\phi(v(\Delta^{\prime}))\Big)$ (III.48)
@@
$\quad\cdot\hat{h}_{s_{N}\left(\Delta^{\prime}\right)}\left[\hat{h}_{s_{N}\left(\Delta^{\prime}\right)}^{-1},\hat{V}_{(v(\Delta^{\prime}),\epsilon)}^{\frac{3}{4}}\right]\Big)\cdot|\beta,\vec{j},\vec{\iota}\rangle\otimes|\phi\rangle$ (III.49)
For the term $T_{\square^{\rho}}^{4}$, we add a characteristic function $\chi_{\square^{\rho}}$ of $\square^{\rho}$ to transpose the integral on $\square^{\rho}$ to $\Sigma$. This allows us to use integration by parts and throw away the boundary term. By using the point-splitting scheme and Eq. (III.31)–Eq. (III.34), this term can be expressed as
@@
$\times\left\{A_{e}^{j}(y),V_{(y,\epsilon)}^{\frac{3}{4}}\right\}\left\{A_{f}^{k}(y),V_{(y,\epsilon)}^{\frac{3}{4}}\right\}$ (III.50)
@@
$\times\Big(\phi\left(f\left(s_{I}\left(\Delta^{\prime\prime}\right)\right)\right)-\phi\left(v\left(\Delta^{\prime\prime}\right)\right)\Big)$ (III.51)
@@
$\quad\cdot\hat{h}_{s_{K}\left(\Delta^{\prime\prime}\right)}\left[\hat{h}_{s_{K}\left(\Delta^{\prime\prime}\right)}^{-1},\hat{V}_{\left(v\left(\Delta^{\prime\prime}\right),\epsilon\right)}^{\frac{3}{4}}\right]\Big)\cdot|\beta,\vec{j},\vec{\iota}\rangle\otimes|\phi\rangle$ (III.52)
@@
$\!\hat{C}_{3}^{\delta}(\!N\!)\!\cdot\!|\beta,\!\vec{j},\!\vec{\iota}\rangle\!\otimes\!|\phi\rangle\!=\!\mp\!\sum_{v}sgn(\!N_{v}\!)\sqrt{\!\sum_{i=1}^{5}\widehat{T^{i,\delta}_{\square^{\rho,v}}}}\!\cdot\!|\beta,\!\vec{j},\!\vec{\iota}\rangle\!\otimes\!|\phi\rangle\!$ (III.53)
@@
$\tilde{\eta}\left(\Psi_{\beta}\right)\left[\Phi_{\beta^{\prime}}\right]:=\sum_{g\in Diff_{np4}/TDiff_{\beta}}\left\langle\hat{U}_{g}\Psi_{\beta}\mid\Phi_{\beta^{\prime}}\right\rangle_{kin}$ (III.54)
@@
$\left<\tilde{\eta}(\Psi_{\beta})|\tilde{\eta}(\Phi_{\beta^{\prime}})\right>_{vtx}:=\tilde{\eta}\left(\Psi_{\beta}\right)\left[\Phi_{\beta^{\prime}}\right]$ (III.55)
By completing $\tilde{\eta}(Cyl_{G})$ using Eq. (III.55), one obtains the vertex Hilbert space $\mathcal{H}_{vtx}$. It turns out that $\hat{C}^{\delta}(N)$ can be promoted to an $\delta$-independent operator on $\mathcal{H}_{vtx}$ via the following dual action:
$[\hat{C}^{*}(N)\cdot\tilde{\eta}(\Psi_{\beta})]\cdot\Phi_{\beta^{\prime}}:=\tilde{\eta}(\Psi_{\beta})(\hat{C}^{\delta}(N)\cdot\Phi_{\beta^{\prime}})$ (III.56)

### v2

@@ III.1 Deparametrization of the Scalar-Tensor Theory
By substituting the function $t(\phi,A,E)$ into Eq. (II.29) and imposing the Hamiltonian constraint $H=0$, one finds that the function $t(\phi,A,E)$ satisfies
$t(\phi,A,E)=\left(\pi+\frac{\tilde{K}_{a}^{i}E_{i}^{a}}{\phi}\right)^{2},$ (III.3)
which implies $t(\phi,A,E)\geq 0$ on the constraint surface. Hence the Hamiltonian constraint can be reformulated as
$C(x)=\pi(x)-(-\frac{\tilde{K}_{a}^{i}E_{i}^{a}}{\phi}\pm\sqrt{t(\phi,A,E)})$ (III.4)
The deparametrized system is composed of the Gaussian constraint (II.26), the diffeomorphism constraint (II.28) and the new Hamiltonian constraint (III.4), which is equivalent to the original constraint system. It remains a first-class constraint system[4].
@@
$\pi\!=\!\pm\!\sqrt{\frac{2p_{ab}p^{ab}}{\phi^{2}}\!-\!\frac{hR}{2}\!-\!\frac{3h}{4\phi^{2}}\!(D_{a}\phi)\!D^{a}\phi\!+\!\frac{h}{\phi}D_{a}D^{a}\phi\!+\!\frac{h\xi(\phi)}{\phi}}$ (III.5)
@@
$-\frac{3h}{4\phi^{2}}(D_{a}\phi)D^{a}\phi+\frac{h}{\phi}D_{a}D^{a}\phi+\frac{h\xi(\phi)}{\phi}$ (III.7)
@@
$C^{\prime}(x)=\pi(x)-(\pm\sqrt{t^{\prime}(\phi,A,E)})$ (III.8)
The deparametrized system is composed of the Gaussian constraint (II.26), the diffeomorphism constraint (II.28), the conformal constraint (II.32) and the new Hamiltonian constraint E(III.8), which is equivalent to the original constraint system.
@@ III.2 Loop Quantization of the Scalar-Tensor Theory
$Cyl^{geo}:=Span\{|\beta,\vec{j},\vec{\iota}\rangle\}$ (III.10)
@@
$\left\langle\psi\mid\psi^{\prime}\right\rangle=\int dg_{1}\ldots dg_{n^{\prime\prime}}\bar{\psi}\left(g_{1},\ldots,g_{n^{\prime\prime}}\right)\psi^{\prime}\left(g_{1},\ldots,g_{n^{\prime\prime}}\right)$ (III.12)
@@
$E(\mathrm{S},\mathrm{g}):=\int_{\mathrm{S}}\epsilon_{abc}E_{i}^{a}\mathrm{g}^{i}$ (III.13)
@@
$\phi:\Sigma\rightarrow\mathbb{R}$ (III.15)
@@
$\langle\phi\mid\phi\rangle=1,\quad\mathrm{and}\quad\left\langle\phi\mid\phi^{\prime}\right\rangle=0\quad\mathrm{whenever}\quad\phi\neq\phi^{\prime}$ (III.17)
@@
$|\phi\rangle[\pi]=\exp\left(-i\int\mathrm{d}^{3}x\phi(x)\pi(x)\right)$ (III.18)
@@
$\hat{\phi}(x)\cdot|\phi\rangle=\phi(x)|\phi\rangle$ (III.19)
@@
$\exp\left(-i\int\mathrm{d}^{3}x\mathrm{f}(x)\hat{\pi}(x)\right)\cdot|\phi\rangle=|\phi+\mathrm{f}\rangle$ (III.20)
@@
$\widehat{\partial_{a}\ldots\partial_{b}\phi}(x)\cdot|\phi\rangle=\partial_{a}\ldots\partial_{b}\phi(x)|\phi\rangle$ (III.21)
@@
$\mathcal{H}_{kin}^{t}=\mathcal{H}_{kin}^{geo}\otimes\mathcal{H}_{kin}^{sf}$ (III.22)
@@
$\hat{U}_{f}\cdot|\beta,\vec{j},\vec{\iota}\rangle\otimes|\phi\rangle=|f(\beta),\vec{j}\circ f^{-1},\vec{\iota}\circ f^{-1}\rangle\otimes|\phi\circ f^{-1}\rangle$ (III.23)
@@
$\hat{P}_{Diff_{\beta}}\cdot|\Psi_{\beta}\rangle:=\frac{1}{n_{\beta}}\sum_{g\in GS_{\beta}}\hat{U}_{g}\cdot|\Psi_{\beta}\rangle$ (III.24)
@@
$\eta\left(\Psi_{\beta}\right)\left[\Phi_{\beta^{\prime}}\right]:=\sum_{f\in Diff/Diff_{\beta}}\left\langle\hat{U}_{f}\hat{P}_{Diff_{\beta}}\Psi_{\beta}\mid\Phi_{\beta^{\prime}}\right\rangle_{kin}$ (III.25)
@@
$\left\langle\eta\left(\Psi_{\beta}\right)\mid\eta\left(\Phi_{\beta^{\prime}}\right)\right\rangle_{Diff}:=\eta\left(\Psi_{\beta}\right)\left[\Phi_{\beta^{\prime}}\right]$ (III.26)
The diffeomorphism invariant Hilbert space $\mathcal{H}_{Diff}$ is defined by the completion of $\eta[Cyl_{G}]$ with respect to the inner product given by Eq. (III.26).
@@ III.3 Quantum Dynamics of the Scalar-Tensor Theory
In the sertor of $\omega(\phi)\neq-\frac{3}{2}$, to quantizethe deparametrized Hamiltonian constraint (III.4), we devide its smearing version into the following three terms:
$C_{1}(N)=\int d^{3}xN(x)\pi(x),$ (III.27)
@@
$\begin{split}\widehat{C}_{1}(N)=&\frac{1}{2i\lambda_{0}}\biggl(\exp\left[i\lambda_{0}\int d^{3}xN(x)\hat{\pi}(x)\right]\\ &-\exp\left[-i\lambda_{0}\int d^{3}xN(x)\hat{\pi}(x)\right]\biggr)\end{split}$ (III.28)
@@
$\times$ $\exp\left(-i\lambda_{0}\int d^{3}y\pi(y)\right)$ (III.29)
@@
$=\frac{-2sgn(\phi(x))}{\lambda_{0}}\left(\sqrt{|\phi(x)|}-\sqrt{|\phi(x)+\lambda_{0}|}\right)|\phi\rangle$ (III.30)
@@
$E_{i}^{a}$ $=\frac{1}{2}\varepsilon^{abc}\varepsilon_{ijk}e_{b}^{j}e_{c}^{k}\operatorname{sgn}\left(\operatorname{det}\left(e_{a}^{i}\right)\right)$ (III.32)
$\operatorname{sgn}\left(\operatorname{det}\left(e_{a}^{i}\right)\right)$ $=\frac{1}{6}\frac{1}{\sqrt{h}}\varepsilon_{ijk}\varepsilon^{abc}e_{a}^{i}e_{b}^{j}e_{c}^{k}$ (III.33)
$e_{a}^{i}$ $=\frac{2}{\gamma}\left\{A_{a}^{i},V\right\}$ (III.34)
$\tilde{K}_{a}^{i}$ $=\gamma^{-\frac{5}{2}}\left\{A_{a}^{i},\left\{H^{E}(1),V\right\}\right\}$ (III.35)
@@
$H^{E}(1)=\frac{1}{2}\int d^{3}xF_{ab}^{j}\frac{\varepsilon_{jkl}E_{k}^{a}E_{l}^{b}}{\sqrt{h}}$ (III.36)
@@
$\times\left\{A_{f}^{n}(y),V_{(y,\epsilon)}^{\frac{2}{3}}\right\}$ (III.37)
@@
$h_{s_{I}}$ $\approx 1-\delta\dot{s}_{I}^{a}A_{a}^{i}\tau_{i}$ (III.38)
$\int d^{3}x\varepsilon^{abc}$ $\approx\sum_{\Delta}\frac{1}{6}\delta^{3}\epsilon\left(s_{I}s_{J}s_{K}\right)\varepsilon^{IJK}\dot{s}_{I}^{a}\dot{s}_{J}^{b}\dot{s}_{K}^{c}$ (III.39)
@@
$\times h_{s_{M^{\prime}}\!\left(\!\Delta^{\prime}\!\right)\!}\!\left\{\!h_{s_{M^{\prime}}\!\left(\!\Delta^{\prime}\!\right)\!}^{-1},V_{\!\left(\!v\left(\Delta^{\prime}\!\right)\!,\epsilon\!\right)\!}^{\frac{2}{3}}\right\}h_{s_{N^{\prime}}\!\left(\!\Delta^{\prime}\!\right)\!}\!\left\{\!h_{s_{N^{\prime}}\!\left(\!\Delta^{\prime}\!\right)\!}^{-1},V_{\!\left(\!v\left(\Delta^{\prime}\!\right)\!,\epsilon\!\right)\!}^{\frac{2}{3}}\!\right\}\!\Big)$ (III.40)
@@
$\quad\left.\cdot\hat{h}_{s_{N}\left(\Delta^{\prime}\right)}\left[\hat{h}_{s_{N}\left(\Delta^{\prime}\right)}^{-1},\hat{V}_{(v(\Delta^{\prime}),\epsilon)}^{\frac{2}{3}}\right]\right)\cdot|\beta,\vec{j},\vec{\iota}\rangle\otimes|\phi\rangle$ (III.41)
@@
$C_{3}(N)=\mp\lim_{\rho\rightarrow 0}\sum_{\square^{\rho}}\operatorname{sgn}\left(N_{\square^{\rho}}\right)\sqrt{T_{\square^{\rho}}}$ (III.42)
@@
$-\left(\int_{\square^{\rho}}d^{3}xN(x)\sqrt{\frac{(3+2\omega(\phi))\xi(\phi)}{\phi}}\sqrt{h}\right)^{2}$ (III.43)
@@
$H^{L}=\frac{1}{2}\frac{\varepsilon_{jmn}\varepsilon_{jkl}\tilde{K}_{a}^{m}\tilde{K}_{b}^{n}E_{k}^{a}E_{l}^{b}}{\sqrt{h}}$ (III.44)
@@
$\cdot\sum_{v\in\beta(v)\cap\square^{\rho}}N(v)(3+2\bar{\omega}(v))\hat{H}^{E}_{v}\cdot|\beta,\vec{j},\vec{\iota}\rangle\otimes|\phi\rangle$ (III.45)
@@
$\cdot\!\sum_{v\in\beta(v)\cap\square^{\rho}}\!N(v)(3\!+\!2\bar{\omega}(v))(\gamma^{2}\!+\!D^{2}(v))\hat{H}^{L}_{v}\cdot|\beta,\vec{j},\vec{\iota}\rangle\otimes|\phi\rangle$ (III.46)
@@
$\cdot\sum_{v(\Delta)=v}\hat{V}_{(v(\Delta),\epsilon)}\cdot|\beta,\vec{j},\vec{\iota}\rangle\otimes|\phi\rangle$ (III.47)
For the term $T_{\square^{\rho}}^{3}$, we use the point-splitting scheme and apply Eq. (III.32)–Eq. (III.35) to express it as
@@
$\times\left\{A_{e}^{m}(y),V_{(y,\epsilon)}^{\frac{3}{4}}\right\}\left\{A_{f}^{n}(y),V_{\left(y,\epsilon\right)}^{\frac{3}{4}}\right\}$ (III.48)
Then, by Eq. (III.38)–Eq. (III.39), it can be regularized as
@@
$\times\Big(\phi(f(s_{L}(\Delta^{\prime})))-\phi(v(\Delta^{\prime}))\Big)$ (III.49)
@@
$\quad\cdot\hat{h}_{s_{N}\left(\Delta^{\prime}\right)}\left[\hat{h}_{s_{N}\left(\Delta^{\prime}\right)}^{-1},\hat{V}_{(v(\Delta^{\prime}),\epsilon)}^{\frac{3}{4}}\right]\Big)\cdot|\beta,\vec{j},\vec{\iota}\rangle\otimes|\phi\rangle$ (III.50)
For the term $T_{\square^{\rho}}^{4}$, we add a characteristic function $\chi_{\square^{\rho}}$ of $\square^{\rho}$ to transpose the integral on $\square^{\rho}$ to $\Sigma$. This allows us to use integration by parts and throw away the boundary term. By using the point-splitting scheme and Eq. (III.32)–Eq. (III.35), this term can be expressed as
@@
$\times\left\{A_{e}^{j}(y),V_{(y,\epsilon)}^{\frac{3}{4}}\right\}\left\{A_{f}^{k}(y),V_{(y,\epsilon)}^{\frac{3}{4}}\right\}$ (III.51)
@@
$\times\Big(\phi\left(f\left(s_{I}\left(\Delta^{\prime\prime}\right)\right)\right)-\phi\left(v\left(\Delta^{\prime\prime}\right)\right)\Big)$ (III.52)
@@
$\quad\cdot\hat{h}_{s_{K}\left(\Delta^{\prime\prime}\right)}\left[\hat{h}_{s_{K}\left(\Delta^{\prime\prime}\right)}^{-1},\hat{V}_{\left(v\left(\Delta^{\prime\prime}\right),\epsilon\right)}^{\frac{3}{4}}\right]\Big)\cdot|\beta,\vec{j},\vec{\iota}\rangle\otimes|\phi\rangle$ (III.53)
@@
$\!\hat{C}_{3}^{\delta}(\!N\!)\!\cdot\!|\beta,\!\vec{j},\!\vec{\iota}\rangle\!\otimes\!|\phi\rangle\!=\!\mp\!\sum_{v}sgn(\!N_{v}\!)\sqrt{\!\sum_{i=1}^{5}\widehat{T^{i,\delta}_{\square^{\rho,v}}}}\!\cdot\!|\beta,\!\vec{j},\!\vec{\iota}\rangle\!\otimes\!|\phi\rangle\!$ (III.54)
@@
$\tilde{\eta}\left(\Psi_{\beta}\right)\left[\Phi_{\beta^{\prime}}\right]:=\sum_{g\in Diff_{np4}/TDiff_{\beta}}\left\langle\hat{U}_{g}\Psi_{\beta}\mid\Phi_{\beta^{\prime}}\right\rangle_{kin}$ (III.55)
@@
$\left<\tilde{\eta}(\Psi_{\beta})|\tilde{\eta}(\Phi_{\beta^{\prime}})\right>_{vtx}:=\tilde{\eta}\left(\Psi_{\beta}\right)\left[\Phi_{\beta^{\prime}}\right]$ (III.56)
By completing $\tilde{\eta}(Cyl_{G})$ using Eq. (III.56), one obtains the vertex Hilbert space $\mathcal{H}_{vtx}$. It turns out that $\hat{C}^{\delta}(N)$ can be promoted to an $\delta$-independent operator on $\mathcal{H}_{vtx}$ via the following dual action:
$[\hat{C}^{*}(N)\cdot\tilde{\eta}(\Psi_{\beta})]\cdot\Phi_{\beta^{\prime}}:=\tilde{\eta}(\Psi_{\beta})(\hat{C}^{\delta}(N)\cdot\Phi_{\beta^{\prime}})$ (III.57)